thomaster (thomaster):

if xy=1 what is the value of: 2(x+y)^2/2 (x-y)^2

jhonyy9 (jhonyy9):

so to solve this exercise you need to know some formules (a+b)^2 = ? (a-b)^2 = ? do you know ?

thomaster (thomaster):

(a+b)^2 = (a+b)(a+b) = a2+ 2ab + b2 (a-b)^2 = (a-b)(a-b) = a2- 2ab + b2

jhonyy9 (jhonyy9):

so than now you just rewrite it using a=x and b=y ok ?

thomaster (thomaster):

yes but if x*y = 1 how do i know the values of x and y it could be 1 * 1 or 2 * 0,5

jhonyy9 (jhonyy9):

no so you know that x*y = 1 from what you can getting that x = 1/y using this you need just substitute inside this equation

jhonyy9 (jhonyy9):

2(x+y)^2/2 (x-y)^2 x = 1/y 2(1/y +y)^2 / so the denominator is 2(x-y)^2 ???

jhonyy9 (jhonyy9):

how is right please ?

thomaster (thomaster):

no idea i'm bad at math xD There are 5 mc anwsers: 1, 2, 4, 16 and 19 If i take x=1 and y=1 it's 8/0 which is impossible if x=2 and y=1/2 the answer is 2,777..... if x=3 and y=1/3 the answer is 1,546..... if x=1/2 and y =2 the answer is 9 if x=1/3 and y=3 the answer is 4 but how can i ever know that x=1/3 and y=3 is the right combination

OpenStudy (anonymous):

I think the equation you wrote is missing an "=" sign.... could you rewrite it using the "equation" button below so we understand correctly

thomaster (thomaster):

already solved it: \[2(x+y)^2/2(x-y)^2 \]\[= 2((x+y)^2 - (x-y)^2) \]\[= 2^{4xy}\]\[= 2^{4*1}\]\[= 16\]

OpenStudy (phi):

There is no unique value for \[ \frac{2(x+y)^2}{2 (x-y)^2} \] do you mean \[ 2(x+y)^2 - 2(x-y)^2 \]?

OpenStudy (phi):

For the 2nd case, expand each square (use FOIL) Example: (x+y)^2= x^2 + 2xy + y^2 and simplify

thomaster (thomaster):

lol i'll just stick with biology and chemistry xD i hate this kind of math :')